While conducting a one-way ANOVA comparing 6 treatments with 10 observations per treatment,  the computed value for SS(Treatment)= 1, and SS(Error)=24. Calculate the value of F.

Round off the answer to 2 decimal digits

A study was conducted to determine the association between the maximum distance at which a highway sign can be read ( in feet) and the age of the driver ( in years). Fourty drivers of various ages were studied. The summary statistics for distance and age are shown below in table :

Mean

Std Deviation

Distance (y)

177.3

6.6

Age(x)

30.04

3.9

Given that the correlation coefficient between distance and age in this sample is r(x,y) = 0.5, calculate the slope of the least-squares regression.

Round off the answer to 2 decimal digits.

The observed frequencies of sales of different colours of cars are shown in the following table:

Category colour

Black

Red

Green

White

Blue

Observed Frequencies

20

30

25

25

47

  

Calculate the chi-square test statistic to test the claim of equal probabilities.

The degree of freedom of goodness of fit-test with 10 categories is equal to……..

Listed below the regression equation of the heights (in inches) and weights (in lb) of 6 randomly selected males(Here y is the weight and x is the height of male)

y with hat on top equals negative 200 plus 5 x space             

Find the predicted weight of a male who is 70.0 inch tall?

A fitted linear regression model is  y with hat on top equals 3 minus x  . The residual at the point (0, 5) is

In contingency tables. When conduction a hypothesis test of independence between variables with two rows and two columns (2x2), if the value of chi square test statistic=58.393, p-value =0.000 and the level of significance =0.01 then reject the null hypotheses of independence between rows and columns.

True

False

  1. In goodness of fit test, for each category, the expected frequeny must be less than 5.

True

False

  1. When conducting a hypothesis test with chi-square analysis, the critical region in a chi-square distribution is always in the left tail.

True

False

Analysis of variance (ANOVA) is a method for testing the hypothesis that three or more population means are equal

True

False

  1. In a one way ANOVA, the value of test statistic F=0.104 and the variance within samples equal 1.23 then the variance between samples is 0.229.

True

False

To test the claim that there is a linear correlation between the two variables x and y with n=5 and r =0.5 then the test statistic t value equals:

1

0.75

0.25

2

The sample data of energy consumptions by households in the Northeast, Midwest and South are given below. Find the variance between the samples? n=5 for each.

Northeast

Midwest

South

Mean( top enclose x )

48.2

35.4

69.8

Variance(S2)

153.2

128.8

183.2

9.97

51.16

1511.46

302.28

The value of the linear correlation coefficient between x and y is 0.4. What is the value of coefficient of determination?

0.2

0.6

-0.4

0.16
Answer #1

due to HomeworkLib policy i am answering starting some question , please repost remaining questions

# slope (b) = r (sy/sx)

=0.5(6.6/3.9)

0.846

NEXT QUESTION,  

# here expected frequency of selection one colour= 1/5 = 20%=0.20

The following table is obtained: Categories Observed Expected (fo-fe)?/fe Black 147*0.20=29.4 Red 147*0.20=29.4 (20-29.4)/29.

# df=n-1 = 9

# LAST QUESTION

coefficient of determination= r​​​​​​2​​​= 0.16

please like ?

please like ?